OpenStudy (anonymous):

Fun Land theme park has a total of 140 males and females who are employed at the park. If the number of males is 20 more than twice the number of females, how many males are employed at the park?

OpenStudy (anonymous):

-x + 2y = -6 -x = -2y -6 x = 2y + 6

OpenStudy (pfenn1):

If x is the number of males employed at the park, then the number of women employed at the park is 140-x. How would you write the expression "number of males is 20 more than twice the number of females" in mathematical terms?